OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence a1=17 and a15=-39? answers to pick. is the answer -143?? a) x=-147 b) x=-143 c) x=-139 d) x=-135

OpenStudy (amistre64):

do you recall the formula for the sum of an arith seq?

OpenStudy (anonymous):

no im just doing makeup work i missed the notes for that day im just rushing trying to get done

OpenStudy (amistre64):

it should be somewhere in your study material ....

OpenStudy (anonymous):

okay could you give to me so i could write down for my records

OpenStudy (amistre64):

\[a_n=a_1+d(n-1)~:~n=1,2,3,...\]

OpenStudy (anonymous):

thank you is there any way you could guide me through the problem?

OpenStudy (amistre64):

use a1 and a15 to find "d" once you know the what d is you replace a15 by a41

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

for the answer i got -143 is that correct?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@hartnn

OpenStudy (amistre64):

a1=17 and a15=-39 -39 = 17 + d(15-1) (-39 -17)/14 = d a41 = 17 + (-39 -17)(40)/14 ... i get -143 as well

OpenStudy (anonymous):

is -143 the answer ? @amistre64